Please help me slove this word problem! What formula do you use to slove this?
Phil wants to make the largest possible rectangular vegetable garden using 18m of fencing. The garden is right behind the back of his house, so he has to fence it only on three sides. Determine the dimensions and the maxium area of the garden.

the anwser should be 9m by 4.5m, 162/4 m^2

Answer by oscargut(686)

You have to find the maximum of area = LW as you know L=18-2W
then you have to find maximum of f(W)=(18-2W)W that is maximum when W=4.5 and L=9
Finally maximum area = 9*4.5=40,5 m^2